OpenStudy (yuii):

Train A leaves Westtown and travels at 50 mph toward Smithville, 330 miles away. At the same time, Train B leaves Smithville and travels at 60 mph toward Westtown. After how many hours do the two trains meet?

OpenStudy (yuii):

3 hours?

OpenStudy (royalranger):

50x + 60x = 330 110x = 330 x = 3

OpenStudy (welshfella):

the trains are approaching each other at a speed of 50 + 60 = 110 mph distance travelled when they meet = 330 time = distance / speed = 330/110 = 3 hours you are correct

OpenStudy (yuii):

Thank you both.

OpenStudy (welshfella):

yw